A stratified simple random sample (STSI) is being planned. Information is available from a pilot study...

  1. A stratified simple random sample (STSI) is being planned. Information is available from a pilot study on the variance s1+k+Cjh1z5OrnWr7nStmfWJyZLRm2wcXoZLSFd4 and cost of data collection Cj2H5lMmRK1A6AB8cwbCkc+NLak8SwiVJWxW6dBN in stratum h. The size of stratum h is known to be MRr7WOXujJFw7GdsQaZw3N+2tEOo8ME509DnNhbg. The cost equation is Ym+LPWfQbYwAAAAASUVORK5CYII=. The total sample size is 2UkGe56P8ozs2VX7rnJ4BfSv6IqqvazUAAAAASUV. Derive the optimal allocation of units (4bgZPkB8ZgfpIzar0yqcAEKZWrI54MpEAAAAASUV) to the strata.
Answer #1

A stratified simple random sample (STSI) is being planned Information is available from a pilot study on the variance Syun ThOur onbiased estimator of Ist = I wn In T is where wn- Nu SRS WOR - vare (5 st) = 3 wir vare (in) Here, var (in) = Suun (Na-nnh h=1 nn h ! N Under SRSDOR, sin is an unbiased estimator of Syun V = Vare (Est) = (Nn-on) Na susun Now, ve n huon – 2 An soGeiven (= e* to men & et_co Wn Syun .cn = c *- co Wn syun Sen = c*_ca hol I wa sogon den put the value of h is og we get nha» C&-lo This ninn n Wnsyon 5 Wh Syun h=1 = n Nn Syun | 5 b= [( ) + & Nh Sy Un hal (Neymans This is of units called (ne optim